find the equation of the line through (-4,6)and perpendicular y=-3x+4
Answer by Alan3354(69443)

find the equation of the line through (-4,6)and perpendicular y=-3x+4
------------
The slope, m, of y=-3x+4 is -3.
The slope of lines perpendicular is the negative inverse, +1/3.
Use
y-y1 = m*(x-x1) where (x1,y1) is the point.
y-6 = (1/3)*(x+4)
3y-18 = x+4
x-3y = -22